OpenStudy (anonymous):

|2x+1|=x-3 aaand the value of x is?

OpenStudy (nubeer):

2x+1 = x-3 x = -4 and -2x-1 = x-3 3x = 2 x=2/3

OpenStudy (anonymous):

@nubeer But if X=-4, and 2/3. |2x+1| = -7 |2x+1| = -2 2/3

OpenStudy (anonymous):

and absolute value cant equal a negative..

hartnn (hartnn):

you are absolutely correct @Mello so this has no solution.

OpenStudy (anonymous):

Oh, I thought I was doing something wrong, ending up with no answer XP Thanks guys

hartnn (hartnn):

welcome :)